Suppose that the typical expenditures for a household equalled 2500 per month, while the cost of purchassing exactly the same items in 2005 was 3000. If 2000 is the base year, the CPI for 2000 equals?
1
Expert's answer
2013-03-25T11:35:59-0400
A consumer price index (CPI) measures changes in the price level of consumer goods and services purchased by households. The CPI in the United States is defined by the Bureau of Labor Statistics as "a measure of the average change over time in the prices paid by urban consumers for a market basket of consumer goods and services."
CPI = total updated cost/base period total cost*100%
If the 2000 year is the base year, then CPI 2000 = 2500/2500*100% = 100%, CPI 2005 = 3000/2500 = 1.2*100% = 120%